A building with a cost of $585,000 has an estimated residual value of $117,000, has an...

  1. A building with a cost of $585,000 has an estimated residual value of $117,000, has an estimated useful life of 36 years, and is depreciated by the straight-line method.

    a. What is the amount of the annual depreciation?
    $

    b. What is the book value at the end of the twentieth year of use?
    $

    c. If at the start of the twenty-first year it is estimated that the remaining life is 20 years and that the residual value is $25,000, what is the depreciation expense for each of the remaining 20 years?
Answer #1

Given:

Cost: $585,000

Residual Value: $117000

Estimated life : 36 years

Method of depreciation : Straight line

a. Amount of annual depreciation

Annual depreciation = (Cost-Residual value ) / Estimated life

=(585000-117000) / 36

=$13,000

Therefore, annual depreciation = $13000

b. Book Value At the end of 20 years :

Cost =585000

Accumulated depreciation for last 20 years = 13000(as calculated above) * 20 years

= 13000*20

=260,000

Book value t the end of the year = cost - Accumulated depreciation

=585000-260000

=325000

Therefore . book value at the end of 20 years =325,000

c. If at the start of the twenty-first year it is estimated that the remaining life is 20 years and that the residual value is $25,000, what is the depreciation expense for each of the remaining 20 years?

Sol :

Book value at the end of 20th year: 325000

Residual value: 25000

Life : 20 years

Annual depreciation = ( book value at the end of 20th year - residual value) / life

= (325000-25000) / 20

= 15000

Annual depreciation from 21st year is $15,000 per year for next 20 years .

NOTE: 1. as there is change in the estimated life and residual value from 21st year , book value at the end of 20th year is taken to calculated depreciation.